OpenStudy (anonymous):

y=3x-24 5x-y=8 Walk me through?

OpenStudy (anonymous):

Always eliminate one of the variables. 3x - y = 24 5x - y = 8 Subtract the two equations. You get: -2x = 16 x = -8 Substitute into one of the equations and find y.

OpenStudy (anonymous):

5x-(3x-24)=8 substitute y from first equation into the second equation 5x-3x+24=8 2x+24=8 2x=-16 x=-8

OpenStudy (anonymous):

oh.